# ZeroDVFS: Zero-Shot LLM-Guided Core and Frequency Allocation for Embedded Platforms

### v1

@@ 5.5 LLM Feature Contribution and Ablation Analysis
5.5.1 Experimental Methodology
We evaluate prediction accuracy across three feature configurations to isolate the contribution of each feature category. The baseline configuration uses only the 17 Tree-sitter syntactic features without benchmark identifiers, representing what can be extracted through conventional static analysis. The augmented configuration adds hardware performance counters collected during execution, capturing runtime behavior. The complete configuration further incorporates all 13 LLM-extracted semantic features.
The evaluation dataset comprises 42 OpenMP benchmarks drawn from the Barcelona OpenMP Tasks Suite (BOTS) and PolybenchC, covering task-parallel algorithms, linear algebra kernels, stencil computations, and data mining workloads. Each benchmark executes across all frequency-core configurations with 8 repetitions, yielding 20,160 samples for Jetson TX2 and 26,880 samples for RubikPi. We employ benchmark-stratified sampling with a 70/10/20 split for training, validation, and testing, ensuring representation of all programs in each partition. The prediction model is a gradient boosting regressor implemented in XGBoost with 100 estimators, maximum depth 6, and learning rate 0.1. These hyperparameters were selected through cross-validation for robustness to feature correlations.
5.5.2 Accuracy on Known Benchmarks
Table 7 presents prediction accuracy across configurations and platforms. The coefficient of determination (R2) measures the proportion of variance explained by the model, while mean absolute percentage error (MAPE) quantifies relative prediction error. The train-test gap indicates potential overfitting.
@@ 5.5.2 Accuracy on Known Benchmarks
Table 7: Top 15 Features by Importance (Claude + TX2 + All Features)
Rank Feature Imp. (%)
1 energy_main_j 28.52
2 context_switches 22.27
3 energy_system_j 16.59
4 instructions 7.15
5 energy_gpu_j 4.80
6 energy_ddr_j 3.23
7 cache_references 2.72
8 energy_denver_j 2.22
9 energy_cpu_j 1.41
10 cache_behavior (LLM) 1.03
11 num_cores 0.75
12 cycles 0.59
13 cache_misses 0.59
14 power_cpu_w 0.57
15 data_dependency (LLM) 0.57
  • Gray rows indicate LLM-extracted features.
@@
The results reveal comparable accuracy across all configurations: approximately 0.94 R2 for Jetson TX2 and 0.97 for RubikPi regardless of whether LLM features are included. Paired t-tests comparing baseline and LLM-augmented configurations yield t=1.67t=-1.67 with $p=0.156$, confirming no statistically significant difference at $\alpha=0.05$. This observation has a clear interpretation.
@@
On random train-test splits where training data contains samples from all 42 benchmarks, the model learns benchmark-specific patterns that transfer to held-out samples of the same programs. In this setting, syntactic features already distinguish between benchmarks sufficiently well; semantic features provide redundant information that the model appropriately ignores. The critical evaluation scenario, namely predicting for entirely new programs, requires a different experimental design; leave-one-out cross-validation remains an avenue for future work to further validate generalization to unseen programs.
5.5.3 Feature Importance Hierarchy
Feature importance analysis using the XGBoost model’s gain metric reveals a clear hierarchy in predictive contributions. Hardware runtime metrics dominate: energy measurements (main CPU, system, GPU, DDR, Denver cores) collectively account for over 50% of total importance, with energy_main_j alone contributing 28.5%. Context switches rank second at 22.3%, followed by instruction counts (7.2%) and cache references (2.7%). This dominance reflects the direct causal relationship between these metrics and execution time, as programs consuming more energy generally run longer.
Table 7 presents the complete feature importance ranking. Among semantic features, cache behavior pattern (1.03%) and data dependency type (0.57%) contribute most significantly, appearing in the top 15 features. While the aggregate contribution of all 13 LLM features totals approximately 5%, this modest importance in the presence of hardware counters does not diminish their value. The key insight is that semantic features encode predictive properties about how code will behave before execution, whereas hardware counters measure observed behavior during specific runs. When hardware counters are unavailable, such as in zero-shot prediction for entirely new programs, semantic features become essential for enabling immediate deployment without exhaustive profiling.
5.5.4 Zero-Shot Generalization
The distinguishing capability enabled by semantic features is zero-shot prediction for programs absent from training data. Traditional approaches using benchmark identifiers achieve high accuracy on known programs but cannot predict for unknown ones; the model has no basis for inference when encountering an unfamiliar identifier. By replacing opaque identifiers with interpretable semantic features, our approach learns relationships between code characteristics and execution time that transfer to novel programs.
The practical deployment workflow for a new program proceeds as follows. Tree-sitter extracts syntactic features in under one second without executing the program. An LLM query retrieves semantic features in under five seconds. The combined feature vector feeds into the trained predictor, yielding an execution time estimate immediately. This entire process completes in seconds without requiring access to target hardware, compared to 8 to 12 hours for exhaustive profiling across all frequency-core configurations.
5.5.5 Cost-Effectiveness Analysis
Table 8 summarizes the practical trade-offs between LLM-based feature extraction and traditional exhaustive profiling. The comparison shows significant cost advantages for the LLM approach across multiple dimensions.
Table 8: LLM Feature Extraction vs. Traditional Profiling
Metric LLM Extraction Traditional
Time per program <<5 seconds 8 to 12 hours
Monetary cost per program \$0.0015 to \$0.018 Hardware + electricity
Requires program execution No Yes
Requires target hardware No Yes
Generalizes to new programs Yes No
Cross-platform applicability Yes Platform-specific
The one-time extraction cost using all three LLMs (\$0.018 per program) enables unlimited subsequent predictions across any number of target platforms. For context, profiling 1,000 programs at \$50/hour labor cost over 8 hours each would exceed \$400,000, making the \$18 total LLM cost for the same workload negligible by comparison. This four-order-of-magnitude cost reduction enables practical deployment of zero-shot prediction in scenarios where exhaustive profiling is economically infeasible.

### v2

@@ 5.5 LLM Feature Contribution and Ablation Analysis
Experimental Methodology. We evaluate three feature configurations: baseline (17 Tree-sitter syntactic), augmented (baseline + 22 hardware counters), complete (augmented + 13 LLM semantic). Dataset: 42 OpenMP benchmarks with 70/10/20 splits. XGBoost with 100 estimators. Details in supplementary material.
Accuracy on Known Benchmarks. Table 6 presents prediction accuracy. R2 measures variance explained, MAPE quantifies relative error, gap indicates overfitting.
@@
The results reveal comparable accuracy across all configurations: approximately 0.94 R2 for Jetson TX2 and 0.97 for RubikPi regardless of whether LLM features are included. Statistical testing confirms no significant difference between baseline and LLM-augmented configurations. This observation has a clear interpretation.
@@
On splits with samples from all benchmarks, syntactic features sufficiently distinguish programs; semantic features provide redundant information. Predicting entirely new programs requires different experimental design; leave-one-out validation remains future work.
Feature Importance Hierarchy. Hardware metrics dominate: energy measurements account for 50%+ importance (energy_main_j 28.5%), context switches 22.3%. LLM features total  5% but encode predictive properties pre-execution vs. observed hardware behavior. When hardware counters unavailable (zero-shot), semantic features enable deployment without profiling.
Zero-Shot Generalization. Semantic features enable zero-shot prediction for unseen programs. Traditional benchmark identifiers fail on unknown programs. Deployment workflow: Tree-sitter extracts syntactic features (under 1s), LLM retrieves semantic features (under 5s), and predictor estimates execution time immediately. This completes in seconds versus 8-12 hours for exhaustive profiling.
Cost-Effectiveness Analysis. LLM extraction completes in under 5 seconds versus 8-12 hours for exhaustive profiling, costs \$0.0015-\$0.018 per program versus substantial hardware costs, requires no target hardware access, and generalizes to new programs without re-profiling. One-time extraction (\$0.018/program) enables unlimited cross-platform predictions. Profiling 1,000 programs at \$50/hour over 8 hours would cost \$400,000 versus \$18 LLM cost (four-order-of-magnitude reduction). Detailed cost-benefit analysis in supplementary material.